OpenStudy (anonymous):

]Choose the equation of the line passing through the point (-4, -2) and perpendicular to y = -x + 6.

OpenStudy (anonymous):

y = -x + 6 y = x - 6 y = -x - 2 y = x + 2

OpenStudy (anonymous):

the equation y=-x+6 is in slope intercept form y=mx+c with slope m=-1 if equation is perpendicular to the given line then slope must be m=1 because the product of the slopes of two perpendicular line is -1 . so the required equation must be y=x+C put x=-4 and y=-2 to find value of C can you do this ?

OpenStudy (anonymous):

yes! thanks!

OpenStudy (anonymous):

yow are welcome :)